OpenStudy (anonymous):

if x1=200cot(θ1) x2=200cot(θ2) x1-x2=200(cot(θ1)-cot(θ2) x1=200cot(14°52/60) x2=200cot(45°10/60) 200(cot(14°52/60)-cot(45°10/60))= 554.58 ft

OpenStudy (anonymous):

what will x1=200cot(θ1) x2=200cot(θ2) x1-x2=200(cot(θ1)-cot(θ2) x1=200cot(18°33/60) x2=200cot(51°33/60) 200(cot(18°33/60)-cot(51°33/60))= ft equal to?

OpenStudy (anonymous):

im so lost :9

OpenStudy (dan815):

you can plug it into calculator

OpenStudy (dan815):

use this one http://www.wolframalpha.com/

OpenStudy (anonymous):

thank you!!